OpenStudy (anonymous):

Divide -3x^3 - 2x^2 - x - 2 by x – 2. A. -3x^2 + 4x + 15, R 32 B. -3x^2 + 4x + 15 C. -3x^2 - 8x - 17 D. -3x^2 - 8x - 17, R –36

OpenStudy (cwrw238):

this division has a remainder since f(2) = -24-8 - 2 - 2 = -36 ( by remainder theorem) so the answer is D

OpenStudy (anonymous):

Ahh, I got it! Thanks

OpenStudy (cwrw238):

yw